OpenStudy (anonymous):

Solve the equation 1 over 3 x + 2 = 10 for x.

OpenStudy (anonymous):

cant you do it

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

3x+2-2 = 10-2 3x = 8 x=8/3

OpenStudy (anonymous):

36 30 24 12

OpenStudy (anonymous):

whtz these

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

Hmm ok

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x +2 } = 10\]